# Learning Heterogeneous Mixture of Scene Experts for Large-scale Neural Radiance Fields

### v1

@@ 4.5 Efficiency Analysis
Efficiency of Switch-NeRF. Table XI compare the efficiency of Switch-NeRF [4] and Mega-NeRF. With only half network parameters, Switch-NeRF attains superior accuracy than Mega-NeRF. Each sub-network in Mega-NeRF employs a foreground and a background NeRF and a set of appearance embeddings (AE), which results in more network parameters. In contrast, Switch-NeRF designs an end-to-end framework which needs only one background NeRF and one set of AE for the whole scene, largely decreasing the parameters while getting better accuracy. Since Switch-NeRF has an extra gating network comparing to Mega-NeRF, it uses slightly more floating point operations (FLOPs) for each point, resulting in slightly more memory and time usage for training. However, Switch-NeRF consumes around 20% less testing memory than Mega-NeRF while only increase minor testing time. These results demonstrate that Switch-NeRF can achieve a much better accuracy with large influencing the efficiency, validating the efficiency and wide applicability of our sparse MoE structure on large-scale scene modeling.
@@
Efficient hash dispatching. We analyze the time and memory efficiency of the hash dispatching of Switch-NeRF++ compared to the optimized dispatching operation in the MoE method [71, 4] in Switch-NeRF. As shown in Table VI, when trained with 100k steps, the MoE dispatching uses 5 more hours and more GPU memory. When rendering images, our method uses only 6.65s to render a $1152\times 864$ image, while the MoE dispatching costs 13.5s. These results show that our efficient dispatching can save time and memory for both training and testing by a large margin compared to the MoE dispatching. The memory is calculated by the Pytorch APIs.
@@
Time and memory overhead for rendering of Switch-NeRF++. In this experiment, we evaluate the time and memory overhead of Switch-NeRF++ for rendering images. As highlighted in Table XIV, our Switch-NeRF++ can render a $1152\times 864$ image with only 6.65s, which is a 16x acceleration compared to Switch-NeRF and a 13x acceleration compared to Mega-NeRF. The memory used by our method is even less than half of these two methods. This experiment clearly verifies the superior efficiency of our method. The efficiency enables our method to be applied to real-world, challenging and large-scale scenes.
@@ 5 Conclusion
We propose Switch-NeRF++, a novel framework for highly scalable Neural Radiance Field (NeRF) modeling. It simultaneously addresses the challenges of leanable decomposition, heterogeneity and efficiency in large-scale NeRF by incorporating the Heterogeneous Mixture of Hash Experts (HMoHE) architecture. Extensive experiments show the superior accuracy and scalability of our method compared to the best-performing methods. Our framework emerges as a practical solution for applying NeRF in real-world large-scale scenes with flexible scalability and high efficiency.

### v2

@@ 4.5 Efficiency Analysis
Efficiency of Switch-NeRF. Table XI compares the efficiency of Switch-NeRF [4] and Mega-NeRF. With only half network parameters, Switch-NeRF attains superior accuracy than Mega-NeRF. Each expert in Mega-NeRF employs a foreground and a background NeRF and a set of appearance embeddings (AE), which results in more network parameters. In contrast, Switch-NeRF designs an end-to-end framework that needs only one background NeRF and one set of AE for the whole scene, largely decreasing the parameters while getting better accuracy. Since Switch-NeRF has an extra gating network compared to Mega-NeRF, it uses slightly more floating point operations (FLOPs) for each point, resulting in slightly more memory and time usage for training. However, Switch-NeRF consumes around 20% less testing memory than Mega-NeRF while only increasing minor testing time. These results demonstrate that Switch-NeRF can achieve a much better accuracy with a large influence on the efficiency, validating the efficiency and wide applicability of our sparse MoE structure on large-scale scene modeling.
@@
Efficient hash dispatching. We analyze the time and memory efficiency of the hash dispatching of Switch-NeRF++ compared to the optimized dispatching in the MoE method [73, 4] in Switch-NeRF. As shown in Table VI, when trained with 100k steps, the MoE dispatching uses 5 more hours and more GPU memory. When rendering images, our method uses only 6.65s to render a $1152\times 864$ image, while the MoE dispatching costs 13.5s. These results show that our efficient dispatching can save time and memory for both training and testing by a large margin compared to the MoE dispatching. The memory is calculated by the Pytorch APIs.
TABLE XVII: Comparison of efficiency with GS-based methods on Building scene. Our method remains competitive in training efficiency on large-scale scenes. While GS-based methods can render efficiently, our model is significantly more compact and parameter-efficient, due to compact hash experts and the sparse MoE network architecture.
Recon. time Rend. time Rend. mem. Model size
Hier-GS 3.20h 0.223s 18147M 855M
CityGaussian 7.25h 0.031s 8985M 787M
Switch-NeRF++ 10.7h 6.650s 2857M 119M
@@
Time and memory overhead for rendering of Switch-NeRF++. In this experiment, we evaluate the time and memory overhead of Switch-NeRF++ for rendering images. As highlighted in Table XI, our Switch-NeRF++ can render a $1152\times 864$ image in only 6.65s, which is a 16x acceleration compared to Switch-NeRF and a 13x acceleration compared to Mega-NeRF. The memory used by our method is even less than half of these two methods. This experiment clearly verifies the superior efficiency of our method. The efficiency enables our method to be applied to real-world, challenging, and large-scale scenes.
Breakdown analysis of efficiency. To better substantiate the efficiency of Switch-NeRF++, we present a detailed breakdown of memory and time usage during training and testing by adding our components to the original Switch-NeRF. The training batch size and steps are aligned with Switch-NeRF++ in our main results in Table I. As shown in Table XVI, using efficient hash gating, dispatching, and hash-based experts leads to substantial improvements. Specifically, replacing large MLP experts with efficient hash experts dramatically reduces memory usage and accelerates both training and inference. This is because most of the computation overhead of Switch-NeRF is in its MLP experts.
Efficiency comparison with GS-based methods. Table XVII reports the comparison of reconstruction time, rendering time, memory usage, and model size on the Building scene with GS-based methods. Our model achieves significantly smaller memory cost and model size, due to its compact hash experts and sparse MoE structure. CityGaussian takes 7.25 hours on 4 GPUs (a coarse stage plus 20 chunks) for training, indicating that Switch-NeRF++ remains competitively efficient in training. While GS-based methods excel in rendering speed, they face scalability and robustness challenges on large-scale scenes with sparse images and noisy point clouds. Switch-NeRF++ provides a more compact, robust, and general solution for large-scale scene modeling.
@@ 5 Conclusion
We propose Switch-NeRF++, a novel framework for highly scalable Neural Radiance Field (NeRF) modeling. It simultaneously addresses the challenges of learnable decomposition, heterogeneity, and efficiency in large-scale NeRF by incorporating the Heterogeneous Mixture of Hash Experts (HMoHE) architecture. Extensive experiments show the superior accuracy and scalability of our method compared to the best-performing methods. Our framework emerges as a practical solution for applying NeRF in real-world large-scale scenes with flexible scalability and high efficiency.